OpenStudy (loser66):

Let A be a real 2 x 2 matrix If A has 2 distinct eigenvalues, then A^2 has 2 distinct eigenvalues, True of False. why? Please, help

OpenStudy (loser66):

@Empty

OpenStudy (loser66):

@phi

OpenStudy (loser66):

@cwrw238

OpenStudy (phi):

I think you can write the matrix as S^-1 L S where L has the eigenvalues on the diagonal, and S is the eigenvectors A^2 = S^-1 L S S^-1 L S= S^-1 L^2 S if the values are distinct say -1 and +1 , after squaring you would get +1 +1 no longer distinct

OpenStudy (loser66):

Since it is False, we can find out a counter example to prove it False What if it is True, how to prove? Like: A is 2 x2 matrix, the determinant of A^2 is nonnegative,. How to prove?

OpenStudy (phi):

Let the eigenvalues be -1 1 make up two eigenvectors and compute A = S^-1 L S

OpenStudy (phi):

example A= -3 -4 2 3

OpenStudy (phi):

eigs(A) ans = 1 -1 eigs(A^2) ans = 1 1

OpenStudy (loser66):

@phi I got you. One more question, if I not lucky enough to pick the right one to manipulate, (I meant if I pick the wrong one which gives me the eigenvalues of A^2 distinct,) I fail the test, right?

OpenStudy (empty):

If the matrix has two eigenvalues then you can diagonalize it: \[A=PDP^{-1}\] where \[D=\left[ \begin{array}c \lambda_1 & 0\\0 & \lambda_2\\\end{array} \right]\] so you can square A to get: \[A^2= (PDP^{-1})( PDP^{-1}) = PD^2P^{-1}\] and D^2 has the form: \[D^2=\left[ \begin{array}c \lambda_1 ^2 & 0\\0 & \lambda_2^2\\\end{array} \right]\] which are the eigenvalues of D, so like @phi is saying, \[\lambda_1 = - \lambda_2 \] would be a counter example. I realizeI probably just wasted my time repeating everything that was just said because I'm a slow typer lol oh well

OpenStudy (phi):

In general, if the values are distinct , then the squares are distinct. So it is a special case that they become "not distinct"

OpenStudy (phi):

@Empty but it looks nicer

OpenStudy (loser66):

Hey!! At the end, it is T or F?

OpenStudy (phi):

It only takes one counter-example to make it false. The statement is false

OpenStudy (loser66):

From @Empty we have lambda1 , lambda 2 (distinct) after then, the eigenvalues of A^2 are lambda1 = - lambda2, (distinct also) ??? I confused

OpenStudy (phi):

He is saying if you start with L1 = - L2 (thus distinct) for matrix A then for matrix A^2 you get L1^2 = L2^2 (no longer distinct)

OpenStudy (loser66):

Got you. Thanks a lot.